INTRACAVITY LASER SPECTROSCOPY OF NiH

\underline{LEAH C. O’BRIEN}, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652; and JAMES J. O’BRIEN, Department of Chemistry, University of Missouri, St Louis, MO 63121-4499.

The visible electronic transitions of NiH have been recorded with rotational resolution by intracavity laser absorption spectroscopy. The gas phase NiH molecules were produced in an electric discharge using a nickel cathode and a trace amount of hydrogen with approx. 1 torr argon. Results of the analysis will be presented.